# MorphSeek: Fine-grained Latent Representation-Level Policy Optimization for Deformable Image Registration

### v1

@@ 1 Introduction
Deformable image registration (DIR) is a highly challenging core task in medical image analysis[0101, 0102, 0103]. Its goal is to establish voxel-wise spatial correspondences between two three-dimensional medical images, thereby enabling precise anatomical alignment. Owing to the pronounced non-rigid, large-scale deformations and inter-subject variability of anatomical structures, DIR is substantially more difficult than generic visual recognition tasks: it must achieve global structural alignment while preserving local geometric consistency[0403zongshuDL2020]. Classical registration methods formulate DIR as a continuous optimization problem and solve for the deformation field via iterative procedures[0201, 0202, 0203, 0204], but their computational cost is extremely high[0301, 0302].
@@
Driven by the rapid progress of deep learning[0401alexnet, 0402attentionisallyouneed], recent approaches adopt end-to-end encoder–decoder architectures to directly map image pairs to deformation fields, achieving significant gains in both efficiency and accuracy[0501voxelmorph, 0502transmorph, 0503swinvoxelmorph, 0504nicetrans, 0505nicenet, 0506pivit, 0507transmatch].
(a) Limited Labeled Data
(b) One-shot Inference
Figure 1: Two Major Challenges Faced by DL-based DIR
@@
Nevertheless, deep learning–based DIR still faces two obstacles: (i) it relies heavily on supervision signals despite extremely limited annotations in most medical scenarios, and (ii) the difficulty of mainstream single-shot inference schemes in handling complex large-deformation cases, which ultimately limits registration accuracy (Figure 1).
@@
The first challenge is to reliably solve high-difficulty, large-deformation registration problems given only a very small number of labeled examples. Complex anatomical structures and large-scale non-rigid deformations often require fine-grained voxel-level supervision to be stably aligned. However, segmentation annotations are exceedingly scarce in most medical settings. As a result, most registration models are forced to rely on unsupervised losses based on image similarity[0701banSZ, 0702ban, 0703banFerrante, 0704banlianhe], whose ability to constrain local boundaries and subtle structures is limited. Existing works mainly strengthen unsupervised registration via pseudo-label generation[0801SPACweibiaoqian, 0806ontheflyweibiaoqian, 0804weibiaoqianSynthMorph], architectural refinements[0802cascade, 0803cascade, 0805cyclemorph, 0807diffVoxelmorph], or new similarity metrics[0808mindmetric, 0809CNNmetric], but comparatively little attention has been paid to maximizing supervision efficiency from a fixed yet very limited set of labels, especially for complex large-deformation cases.
@@
The second challenge stems from the fact that most deep learning–based DIR models perform inference via a single forward pass, i.e., they predict the deformation field in one shot[0501voxelmorph, 0502transmorph, 0807diffVoxelmorph]. In scenarios involving large-scale non-rigid deformations, such as thoracic or abdominal registration, such models often fit only the global structural differences while struggling to reliably recover local boundaries and fine geometric details. To improve alignment under complex deformations, several methods introduce step-wise registration[0901VRNetfine, 0902fine, 0903fine, 0904RIIRstep, 0905LapIRN], decomposing a large deformation into a sequence of incremental updates to realize coarse-to-fine optimization. However, existing step-wise frameworks typically rely on manually designed, fixed cascaded structures and lack a learnable multi-step decision policy. Reinforcement learning (RL) has been explored for image registration because its stochastic, Markov decision process is naturally compatible with step-wise optimization. However, most existing RL-based registration methods are confined to low-dimensional rigid transformations[1101RL_G, 1102RL_G, 1103RL_G, 1104RL_G, 1105RL_G, 1106RL_G]. Directly treating a full 3D deformable field as the action space would make memory consumption and sampling cost prohibitive, severely limiting the applicability of RL to real-world deformable registration.
@@
To address these issues, we propose MorphSeek, which reformulates deformable registration as latent-space policy optimization by introducing a sampleable high-resolution latent representation at the top encoder layer and treating it as the policy action, thereby avoiding RL reasoning directly in the million-dimensional deformation field while preserving fine spatial granularity. The framework first performs unsupervised warm-up to shape a stable latent space and then applies Group Relative Policy Optimization (GRPO) with multiple trajectories and multiple steps under weak supervision, repeatedly reusing scarce labels for coarse-to-fine refinement. To make such high-dimensional policies trainable, we further propose Latent-Dimension Variance Normalization (LDVN), which statistically controls the variance of log-likelihoods and yields stable, unbiased gradients for scalable 3D dense prediction.
@@
  • We propose LDVN, the first general statistical normalization scheme that stabilizes GRPO in high-dimensional dense prediction settings. We show that LDVN controls the variance of the log-likelihood without altering the gradient direction or introducing bias, allowing GRPO to operate stably in high-dimensional 3D feature spaces and providing both theoretical and practical support for applying RL to dense prediction tasks.
@@ 2.1 DL-Based Deformable Medical Image Registration
Early DL-based DIR methods were fully supervised using deformation vector fields (DVFs)[0601quanRGS, 0602quanCTS, 0603quanCTS, 0604quanRGS_RL, 0605quanRGS, 0606quanCTS]. After Hu et al. proposed using anatomical segmentations instead of DVFs for supervision[0701banSZ], such fully supervised strategies became less common. Since VoxelMorph[0501voxelmorph], a U-Net-style CNN trainable in an unsupervised manner, subsequent studies have largely evolved within this unsupervised U-Net-style paradigm[0504nicetrans, 0807diffVoxelmorph, 0505nicenet, 0507transmatch, 0502transmorph, 0503swinvoxelmorph].
Meanwhile, leveraging segmentation labels to further improve registration has become an active research direction. Hu et al. extended the label-driven idea and systematically discussed the advantages of segmentation-supervised training over purely unsupervised objectives[0702ban]. Ferrante et al. used segmentation labels to guide the weighting of different similarity terms during registration[0703banFerrante]. Zhou et al. proposed macJNet[0704banlianhe], which jointly learns two segmentation networks and one registration network.
There have also been attempts to combine unsupervised and weakly-supervised learning. Li et al. combined segmentation-labeled and unlabeled image pairs for registration using consistency regularization in a student-teacher framework[1201semi]. Unsupervised models such as VoxelMorph[0501voxelmorph] often include hybrid objectives that combine image-similarity and label-based losses. Chen et al. proposed a training strategy that first performs unsupervised pretraining with randomly generated images and then fine-tunes on the target task, maintaining strong performance when domain-specific data are limited[1202]. However, these approaches still do not simultaneously exploit both unlabeled and labeled data from the same domain to maximally optimize the registration model.
From another perspective, coarse-to-fine registration has been extensively explored. ULAE-Net[1301ulaenet] performs step-wise registration by repeatedly applying the network. LapIRN[0905LapIRN] cascades multiple Laplacian pyramid networks to implement coarse-to-fine alignment. However, these methods use fixed, deterministic schedules and lack adaptive exploration of optimal registration strategies.
@@ 2.2 Reinforcement Learning in DIR
In recent years, reinforcement learning (RL) has advanced rapidly in decision-making, robotics, and large-model reasoning[1002, 1001deepseek]. However, when applied to dense prediction tasks, the continuous and high-dimensional state and action spaces—particularly in 3D—lead to unstable training, high exploration cost, and substantial memory overhead; this bottleneck is especially pronounced in deformable image registration (DIR).
Krebs et al. proposed an agent-based non-rigid registration framework that reduces the DIR action space from dense DVFs to a statistical deformation model (PCA over B-spline–parameterized deformations), significantly lowering the action dimensionality[0604quanRGS_RL]. However, their method requires dense DVFs as supervision, which is impractical in contemporary settings. Luo et al. introduced SPAC[0801SPACweibiaoqian], which compresses a pair into a 64-D plan and relies on an extra critic for stability (SAC-based), which complicates deployment. Moreover, the 64-D bottleneck discards spatial detail, limiting performance.
@@ 3.3 Multi-Trajectory GRPO for Step-Wise Registration
After warm-up pretraining, we fine-tune the encoder–decoder with segmentation labels under the GRPO framework to further improve registration accuracy. In this stage, the encoder’s stochastic output distribution is treated as a policy, denoted by $\pi(\mathbf{z}\mid\mathbf{f}_{L})$, where the state $s_{t}$ is the current registration pair $\{I_{m}^{t-1},I_{f}\}$ and the action $a_{t}$ is the sampled latent $\mathbf{z}$. At each fine-tuning step $t$, we generate a group of trajectories per sample to enable exploration through encoder stochasticity.
@@
where $\bar{R}$ and $\sigma_{R}$ are the mean and standard deviation of $\{R^{(j)}\}_{j=1}^{J}$ for the current sample, and $\epsilon=10^{-8}$ prevents division by zero. We also compute the relative log-likelihood within the group:
@@
where $\boldsymbol{\mu},\boldsymbol{\sigma}$ are the encoder outputs for the current state (shared across the $J$ samples), and $\overline{\log\pi}$ is the group mean of $\log\pi(\mathbf{z}^{(j)}\mid\boldsymbol{\mu},\boldsymbol{\sigma})$.
However, for conventional backbones[0502transmorph, 0504nicetrans], we observe that the latent dimensionality $\mathbf{z}$ (denoted as $N=C_{L}\times H_{L}\times W_{L}\times D_{L}$) often reaches tens of thousands, which is far larger than in typical GRPO applications. If we directly sum over all $N$ positions when computing $\log\pi$, the variance of $\log\pi$ within each group will escalate rapidly with $N$. This leads to impaired exploration discriminability and unstable training.
To address this issue, we propose Latent-Dimension Variance Normalization (LDVN), which introduces a scaling factor $s$ in the computation of $\log\pi$:
$\begin{split}\log\pi&(\mathbf{z}|\boldsymbol{\mu},\boldsymbol{\sigma})\\ &=-\frac{1}{2s}\sum_{i=1}^{N}\left[\left(\frac{z_{i}-\mu_{i}}{\tau\sigma_{i}}\right)^{2}+\log(2\pi\tau^{2}\sigma_{i}^{2})\right].\end{split}$ (12)
For different image resolutions and backbone architectures, we dynamically adjust $s$ to constrain $\mathrm{std}(\log\tilde{\pi}^{(j)})$ to a constant value. In doing so, we ensure statistical consistency with Eq. 10 and 11, thereby achieving stable policy updates while preserving fine-grained stochasticity.
We can prove the unbiasedness of LDVN through a simple derivation. For any $s>0$, $b\in\mathbb{R}$,
$(\frac{1}{s}\log\pi^{(j)}+b)-\overline{(\frac{1}{s}\log\pi+b)}=\frac{1}{s}(\log\pi^{(j)}-\overline{\log\pi}).$ (13)
Hence LDVN only rescales within-group relative log-likelihoods without changing their ordering or the gradient direction; when multiplied by the zero-mean advantage in Eq. 10, it is equivalent to a step-size rescaling, leaving the estimator unbiased. Note that LDVN is only applied to the statistics of $\log\pi$ in the policy loss, and does not participate in the reparameterization sampling of Eq. 5, thus it does not alter the sampling distribution of $\pi(\mathbf{z}|\boldsymbol{\mu},\boldsymbol{\sigma})$ or the exploration temperature $\tau$.
@@
   Update $\boldsymbol{\boldsymbol{\theta}}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{grpo}}$
   Update $\Phi_{t},I_{m}^{t}$ via Eqs. 17 and 18
@@
Regarding the specific value of $s$, we denote the summation terms in Eq. 12 as $X_{i}$. It can be proved that a. with bounded second moments and weak dependence, $\mathrm{Var}(\sum_{i=1}^{N}X_{i})=O(N)$, and b. choosing $s\propto\sqrt{N}$ keeps $\mathrm{Var}(\frac{1}{s}\sum_{i=1}^{N}X_{i})=O(1)$ and $\mathrm{std}(\log\tilde{\pi}^{(j)})=O(1)$. Therefore, the value of $s$ is typically set to $\sqrt{N}$. (Detailed derivation is provided in the Appendix.)
@@
$\mathcal{L}_{\text{Dice}}(\boldsymbol{\theta})\;=\;\frac{1}{J}\,\sum_{j=1}^{J}\,\Big[\,1-\mathrm{Dice}\!\big(S_{f},\,S_{m\circ\Phi_{t}^{(j)}}\big)\Big].$ (15)
Note that $\mathcal{L}_{\text{Dice}}$ is computed with soft labels via trilinear interpolation to ensure differentiability, whereas the reward in $\mathcal{L}_{\text{policy}}$ does not backpropagate gradients and is computed with hard labels to more faithfully reflect the task metric.
Figure 3: The Performance of MorphSeek Across Three Different Tasks. Labels are overlaid only for the two abdominal datasets; OASIS is left unlabeled to avoid clutter from its 35 foreground classes.
To prevent catastrophic forgetting of the representations learned during warm-up and to maintain smooth and physically plausible deformations, we retain the warm-up objective as a regularizer. It is computed through Eqs. 7 and 8 rather than sample averaging. The overall loss for GRPO fine-tuning is [^1] The encoder parameterizes a Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\sigma}^{2})$ in both stages. During warm-up, we regularize it toward $\mathcal{N}(\mathbf{0},\mathbf{I})$ via the KL term in Eq. 8. In GRPO fine-tuning, this same KL divergence (evaluated with $\tau=0$ as in warm-up) is retained to maintain a fixed-prior trust region.

### v3

@@ 1 Introduction
Deformable image registration (DIR) is a highly challenging core task in medical image analysis[31, 46, 40]. Its goal is to establish voxel-wise spatial correspondences between two three-dimensional medical images, thereby enabling precise anatomical alignment. Owing to the pronounced non-rigid, large-scale deformations and inter-subject variability of anatomical structures, DIR is substantially more difficult than generic visual recognition tasks: it must achieve global structural alignment while preserving local geometric consistency[12]. Classical registration methods formulate DIR as a continuous optimization problem and solve for the deformation field via iterative procedures[1, 2, 42, 48], but their computational cost is extremely high[38, 32].
@@
Driven by the rapid progress of deep learning[23, 49], recent approaches adopt end-to-end encoder–decoder architectures to directly map image pairs to deformation fields, achieving significant gains in both efficiency and accuracy[3, 5, 59, 35, 34, 30, 7].
@@
Nevertheless, deep learning–based DIR still faces two obstacles: (i) it relies heavily on supervision signals despite extremely limited annotations in most medical scenarios, and (ii) mainstream single-shot inference schemes remain challenged by complex large deformations, which ultimately limits registration accuracy.
@@
The first challenge is to reliably solve high-difficulty, large-deformation registration problems given only a very small number of labeled examples. Complex anatomical structures and large-scale non-rigid deformations often require fine-grained voxel-level supervision to be stably aligned. However, segmentation annotations are exceedingly scarce in most medical settings. As a result, most registration models are forced to rely on unsupervised losses based on image similarity[18, 19, 11, 57], whose ability to constrain local boundaries and subtle structures is limited. Existing works mainly strengthen unsupervised registration via pseudo-label generation[28, 51, 16], architectural refinements[54, 55, 21, 8], or new similarity metrics[14, 13], but comparatively little attention has been paid to maximizing supervision efficiency from a fixed yet very limited set of labels, especially for complex large-deformation cases.
@@
The second challenge stems from the fact that most deep learning–based DIR models perform inference via a single forward pass, i.e., they predict the deformation field in one shot[3, 5, 8]. In scenarios involving large-scale non-rigid deformations, such as thoracic or abdominal registration, such models often fit only the global structural differences while struggling to reliably recover local boundaries and fine geometric details. To improve alignment under complex deformations, several methods introduce step-wise registration[20, 56, 58, 53, 39], decomposing a large deformation into a sequence of incremental updates to realize coarse-to-fine optimization. However, existing step-wise frameworks typically rely on manually designed, fixed cascaded structures and lack a learnable multi-step decision policy. Reinforcement learning (RL) has been explored for image registration because its stochastic, Markov decision process is naturally compatible with step-wise optimization. However, most existing RL-based registration methods are confined to low-dimensional rigid transformations[26, 27, 29, 37, 17, 47]. Directly treating a full 3D deformable field as the action space would make memory consumption and sampling cost prohibitive, severely limiting the applicability of RL to real-world deformable registration.
@@
To address these issues, we propose MorphSeek, which reformulates deformable registration as latent-space policy optimization by introducing a sampleable high-resolution latent representation at the top encoder layer and treating it as the policy action, thereby avoiding RL reasoning directly in the million-dimensional deformation field while preserving fine spatial granularity. The framework first performs unsupervised warm-up to shape a stable latent space and then applies Group Relative Policy Optimization (GRPO) with multiple trajectories and multiple steps under weak supervision, repeatedly reusing scarce labels for coarse-to-fine refinement. To make such high-dimensional policies trainable, we further propose Latent-Dimension Variance Normalization (LDVN), which controls the variance of log-likelihoods and provides direction-preserving, scale-controlled policy updates for scalable 3D dense prediction.
@@
  • We propose LDVN, a statistical normalization scheme that stabilizes GRPO in high-dimensional dense prediction settings. We show that LDVN controls the variance of the log-likelihood while preserving policy-gradient direction, allowing GRPO to operate stably in high-dimensional 3D feature spaces and providing both theoretical and practical support for applying RL to dense prediction tasks.
@@ 2.1 DL-Based Deformable Medical Image Registration
Early DL-based DIR methods were fully supervised using deformation vector fields (DVFs)[44, 52, 50, 22, 45, 4]. After Hu et al. proposed using anatomical segmentations instead of DVFs for supervision[18], such fully supervised strategies became less common. Since VoxelMorph[3], a U-Net-style CNN trainable in an unsupervised manner, subsequent studies have largely evolved within this unsupervised U-Net-style paradigm[35, 8, 34, 7, 5, 59].
Meanwhile, leveraging segmentation labels to further improve registration has become an active research direction. Hu et al. extended the label-driven idea and systematically discussed the advantages of segmentation-supervised training over purely unsupervised objectives[19]. Ferrante et al. used segmentation labels to guide the weighting of different similarity terms during registration[11]. Zhou et al. proposed macJNet[57], which jointly learns two segmentation networks and one registration network.
There have also been attempts to combine unsupervised and weakly-supervised learning. Li et al. combined segmentation-labeled and unlabeled image pairs for registration using consistency regularization in a student-teacher framework[25]. Unsupervised models such as VoxelMorph[3] often include hybrid objectives that combine image-similarity and label-based losses. Chen et al. proposed a training strategy that first performs unsupervised pretraining with randomly generated images and then fine-tunes on the target task, maintaining strong performance when domain-specific data are limited[6]. However, these approaches still do not simultaneously exploit both unlabeled and labeled data from the same domain to maximally optimize the registration model.
From another perspective, coarse-to-fine registration has been extensively explored. ULAE-Net[43] performs step-wise registration by repeatedly applying the network. LapIRN[39] cascades multiple Laplacian pyramid networks to implement coarse-to-fine alignment. However, these methods use fixed, deterministic schedules and lack adaptive exploration of optimal registration strategies.
@@ 2.2 Reinforcement Learning in DIR
In recent years, reinforcement learning (RL) has advanced rapidly in decision-making, robotics, and large-model reasoning[41, 9]. However, when applied to dense prediction tasks, the continuous and high-dimensional state and action spaces—particularly in 3D—lead to unstable training, high exploration cost, and substantial memory overhead; this bottleneck is especially pronounced in deformable image registration (DIR).
Krebs et al. proposed an agent-based non-rigid registration framework that reduces the DIR action space from dense DVFs to a statistical deformation model (PCA over B-spline–parameterized deformations), significantly lowering the action dimensionality[22]. However, their method requires dense DVFs as supervision, which is impractical in contemporary settings. Luo et al. introduced SPAC[28], which compresses a pair into a 64-D plan and relies on an extra critic for stability (SAC-based), which complicates deployment. Moreover, the 64-D bottleneck discards spatial detail, limiting performance.
@@ 3.2 Warm-up Priors for Stable Policy Optimization
This deterministic warm-up forces anatomical information into the mean code before policy sampling and empirically reduces posterior-collapse risk, preserving the stochastic output variance later required by GRPO exploration.
@@ 3.3 Multi-Trajectory GRPO for Step-Wise Registration
After warm-up pretraining, we fine-tune the encoder–decoder with segmentation labels under the GRPO framework to further improve registration accuracy. In this stage, the encoder’s stochastic output distribution is treated as a latent policy, parameterized as $\pi(\mathbf{z}\mid\boldsymbol{\mu},\boldsymbol{\sigma})$ from the current state feature $\mathbf{f}_{L}$, where the state $s_{t}$ is the current registration pair $\{I_{m}^{t-1},I_{f}\}$ and the action $a_{t}$ is the sampled latent $\mathbf{z}$. Here, $t$ denotes the refinement step within one forward pass, and we initialize the cumulative deformation as $\Phi_{0}=\mathrm{Id}$. At each fine-tuning step $t$, we generate a group of trajectories per sample to enable exploration through encoder stochasticity.
@@
where $\bar{R}$ and $\sigma_{R}$ are the mean and standard deviation of $\{R^{(j)}\}_{j=1}^{J}$ for the current sample, and $\epsilon=10^{-8}$ prevents division by zero. This per-sample normalization also acts as an implicit hard-case reweighting: without it, easy cases with large absolute gains would dominate the gradients, whereas standardized advantages preserve learning signal from anatomically difficult pairs. We also compute the relative log-likelihood within the group:
@@
where $\boldsymbol{\mu},\boldsymbol{\sigma}$ are the policy parameters produced from the current state feature $\mathbf{f}_{L}$ (shared across the $J$ samples), and $\overline{\log\pi}$ is the group mean of $\log\pi(\mathbf{z}^{(j)}\mid\boldsymbol{\mu},\boldsymbol{\sigma})$.
However, for conventional backbones[5, 35], the latent dimensionality $N=C_{L}\times H_{L}\times W_{L}\times D_{L}$ often reaches tens of thousands, which is far larger than in typical GRPO applications. Directly summing over all latent dimensions can make within-group relative log-likelihoods numerically unstable, weakening exploration discrimination and destabilizing training.
We therefore introduce Latent-Dimension Variance Normalization (LDVN), which rescales the log-likelihood as
$\begin{split}\log\pi&(\mathbf{z}\mid\boldsymbol{\mu},\boldsymbol{\sigma})\\ &=-\frac{1}{2s}\sum_{i=1}^{N}\left[\left(\frac{z_{i}-\mu_{i}}{\tau\sigma_{i}}\right)^{2}+\log(2\pi\tau^{2}\sigma_{i}^{2})\right].\end{split}$ (12)
where $s$ controls the scale of within-group relative log-likelihoods across different latent dimensionalities. In practice, setting $s\propto\sqrt{N}$ keeps GRPO updates numerically stable while preserving within-group ordering and the policy-gradient direction. LDVN only affects the policy-loss statistics and does not alter the sampling distribution of $\pi(\mathbf{z}\mid\boldsymbol{\mu},\boldsymbol{\sigma})$ or the exploration temperature $\tau$. Detailed variance analysis and empirical verification are provided in the supplement.
@@
   Update $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{grpo}}$
   Update $\Phi_{t},I_{m}^{t}$ via Eqs. 16 and 17
@@
In practice, we set $s=\sqrt{N}$ by default and defer the variance derivation, weak-dependence analysis, and ablations to the supplement.
@@
$\mathcal{L}_{\text{Dice}}(\boldsymbol{\theta})\;=\;\frac{1}{J}\,\sum_{j=1}^{J}\,\Big[\,1-\mathrm{Dice}\!\big(S_{f},\,S_{m\circ\Phi^{(j)}_{t}}\big)\Big$]. (14)
Note that $\mathcal{L}_{\text{Dice}}$ is computed with soft labels via trilinear interpolation to ensure differentiability, whereas the reward in $\mathcal{L}_{\text{policy}}$ does not backpropagate gradients and is computed with hard labels to more faithfully reflect the task metric. While $\mathcal{L}_{\text{Dice}}$ supplies one deterministic overlap signal, $\mathcal{L}_{\text{policy}}$ ranks $J$ sampled hypotheses at each of $T$ refinement steps for the same labeled pair. Each pair therefore yields $T\times J$ relative supervision events, which helps explain MorphSeek’s label-efficiency gains.
To prevent catastrophic forgetting of the representations learned during warm-up and to maintain smooth and physically plausible deformations, we retain the warm-up objective as a regularizer during GRPO. Beyond stabilization, $\mathcal{L}_{\text{warm}}$ acts as an anatomy-preserving prior that keeps optimization close to the warm-up manifold, reducing reward hacking and implausible but numerically favorable deformations. It is computed through Eqs. 7 and 8 rather than sample averaging. The overall loss for GRPO fine-tuning is [^1] The encoder parameterizes a Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\sigma}^{2})$ in both stages. During warm-up, we regularize it toward $\mathcal{N}(\mathbf{0},\mathbf{I})$ via the KL term in Eq. 8. In GRPO fine-tuning, this same KL divergence (evaluated with $\tau=0$ as in warm-up) is retained to maintain a fixed-prior trust region.